OpenStudy (anonymous):

If the length of each side of a cube is halved, is the surface area halved? Use an example to support your answer.

OpenStudy (anonymous):

Please help! :)

OpenStudy (anonymous):

nope it wouldn't be halved but rather be one fourth of it; The surface area of a cube = 6a^2 so lets use as the first cube which would be the bigger one that 'a' the side = 4; and the second one a = 2. So for the first cube; SA = 6a^2 SA =6(4)^2 SA =6(16) SA =96 Now for the second one; SA = 6a^2 SA =6(2)^2 SA =6(4) SA =24 so lets divide them, 96/24 = 4 so the bigger one is four times the smaller or 24/96 = 1/4 so it represents that the smaller is one fourth of the bigger one. Another way to solve this is by using the formula; a^2 4^2 a^2 2^2 --- = --- --- = --- b^2 2^2 b^2 4^2 a^2 16 a^2 4 --- = --- --- = --- b^2 4 b^2 16 a^2 4 a^2 1 --- = --- = 4 --- = --- b^2 1 b^2 4 Again, the four would be if your saying that the bigger one is four times the smaller, and the 1/4 is used for saying that the smaller is one fourth the big one.